# BioVL-QR: Egocentric Biochemical Vision-and-Language Dataset Using Micro QR Codes

### v2

@@ 1 Introduction
Scientific experiments contain multimodality in nature. In real-world scenarios, a typical approach to reproducing experiments allows researchers to read protocols and execute the experiments. Our overarching goal is to construct multimodal systems to help them reproduce experiments accurately. For example, it is useful to have a system that detects human errors and verbalizes how to recover them to prevent researchers from failing to reproduce experiments. To achieve this goal, vision-and-language technologies, such as video-text alignment [2], are effective approaches. Video-text alignment allows researchers to visually verify the correct procedure.
@@
Previously, a few researchers attempted to focus on the vision-and-language tasks in the field of biochemistry [3, 4, 5]. The key challenge in improving reproducibility is detecting equipment, reagents, and containers because the lab environment is scattered by filling objects on the table, and some objects are even indistinguishable. For example, it is difficult to distinguish tubes that look identical but contain different reagents. In Fig. 1, distinguishing between “GP1” and “GP2” is challenging due to their identical appearance. However, correctly identifying them is crucial for real-world applications, as their chemical properties differ significantly. Previous studies [2, 3] assumed that such visually-indistinguishable objects were manually annotated and given for downstream tasks, but this is costly.
To tackle this, we focus on Micro QR Codes [6] to detect objects automatically. In this paper, we introduce BioVL-QR, a biochemical vision-and-language dataset containing 23 egocentric experiment videos, protocols, and alignment annotation, where Micro QR Codes are attached (Fig. 1). In BioVL-QR, the objects appearing in the videos are attached with Micro QR Codes. In our preliminary studies, we found that detecting objects only using Micro QR Codes is still difficult because the researchers manipulate objects, frequently causing blur and occlusion. To address this, we also propose a novel object labeling method by combining a Micro QR Code detector and an off-the-shelf hand object detector [7]. Based on BioVL-QR, we conducted a task of step localization [2, 4, 8], localizing the steps described in procedural texts within an instructional video. The experimental results show that our approach improves the understanding of biochemical videos.
@@ 2 Related work
Egocentric vision-and-language videos are important for advancements in fields such as AR [9] and robotics [10]. Large-scale egocentric vision-and-language datasets are actively being created [11, 12]. These large-scale datasets do not include biochemistry videos primarily due to the difficulty of domain-specific annotation. Specifically, annotating the experimental procedures and labeling the objects in the videos require a high level of expertise in the relevant scientific domain. This makes the annotation process both time-consuming and resource-intensive.

### v3

@@ 1 Introduction
Scientific experiments contain multimodality in nature. In real-world scenarios, a typical approach to reproducing experiments allows researchers to read protocols and execute the experiments. However, this process is prone to human errors. Therefore, constructing multimodal systems that support experimental execution can be a promising way to enhance reproducibility. For example, it is useful to have a system that detects human errors and verbalizes how to recover them to prevent researchers from failing to reproduce experiments. To construct systems of this kind, vision-and-language technologies, such as step localization [2, 3, 4], are effective approaches. Step localization allows researchers to visually verify the correct procedure.
@@
Previously, a few researchers attempted to focus on the vision-and-language tasks in the field of biochemistry [3, 5, 6]. The key challenge in such tasks is detecting equipment, reagents, and containers because the lab environment is scattered by filling objects on the table, and some objects are even indistinguishable. For example, it is difficult to distinguish tubes that look identical but contain different reagents. In Fig. 1, distinguishing between “GP1” and “GP2” is challenging due to their identical appearance. However, correctly identifying them is crucial for real-world applications, as their chemical properties differ significantly. Previous studies [2, 5] assumed that such visually-indistinguishable objects were manually annotated and given for downstream tasks, but this is costly.
To tackle this, we focus on Micro QR Codes [7] to detect objects automatically. In this paper, we introduce BioVL-QR, a biochemical vision-and-language dataset containing 23 egocentric experiment videos, protocols, and alignment annotation, where Micro QR Codes are attached (Fig. 1). In BioVL-QR, the objects appearing in the videos are attached with Micro QR Codes. In our preliminary studies, we found that detecting objects only using Micro QR Codes is still difficult because the researchers manipulate objects, frequently causing blur and occlusion. To address this, we also propose a novel object labeling method by combining a Micro QR Code detector and an off-the-shelf hand object detector [8]. Based on BioVL-QR, we conducted a task of step localization, localizing the steps described in procedural texts within an instructional video. The experimental results show that our approach improves the understanding of biochemical videos.
In summary, our main contributions are: (1) the BioVL-QR dataset, (2) an object labeling method, and (3) baseline methods for step localization that use Micro QR codes and the proposed object labeling method.
@@ 2 Related work
Egocentric vision-and-language datasets are important for advancements in fields such as AR [9] and robotics [10]. Large-scale egocentric vision-and-language datasets are actively being created [11, 12]. These large-scale datasets do not include biochemistry videos primarily due to the difficulty of domain-specific annotation. Specifically, annotating the experimental procedures and labeling the objects in the videos require a high level of expertise in the relevant scientific domain. This makes the annotation process both time-consuming and resource-intensive.
@@ 6 Conclusion
Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers 25K21274 and Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures (JHPCN) in Japan (Project ID: jh241007).